# Few-Shot Learning from Gigapixel Images via Hierarchical Vision-Language Alignment and Modeling

### v4

@@ 1 Introduction
Whole slide image (WSI) classification is a central task in computational pathology (CPath), enabling cancer diagnosis, subtyping, and prognosis prediction [50, 51, 10]. With gigapixel resolution, WSIs contain detailed spatial information ranging from coarse tissue structures to fine-grained cellular morphology [20], which is crucial for accurate interpretation in these diagnostic tasks. To handle their large size and the absence of fine-grained annotations, multiple instance learning (MIL) is widely adopted [8, 28]. In MIL, each WSI is treated as a bag of instances (patches), and a slide-level label is predicted through a feature aggregator with supervision only at the slide level (Figure 1(A)). However, traditional MIL models [27, 33, 38, 47, 54, 34] face key challenges in real-world clinical settings [23]: (1) they rely on large labeled datasets, which are difficult to obtain due to privacy concerns [40] and the rarity of certain diseases [32], making them ineffective in few-shot scenarios where labeled WSIs are scarce [44, 48]; (2) they use only visual features, making them highly sensitive to staining variability [36, 11] and domain shifts [37, 48]. These challenges call for a data-efficient approach that leverages prior domain knowledge for robust learning under limited supervision.
@@
Vision-language models (VLMs) such as CLIP [46], BLIP [35], and Flamingo [2] have demonstrated strong zero- and few-shot transfer performance by learning joint image-text embeddings through contrastive learning [30, 55, 49, 12]. However, their training in natural images and generic captions limits their effectiveness in CPath, where domain-specific semantics are critical. To address this issue, domain-adapted VLMs such as PLIP [25], QuiltNet [26], and CONCH [39] leverage large-scale pathology patch-text datasets and improve robustness and generalization at the patch level. Motivated by these capabilities, recent efforts [44, 45, 19] have integrated them into MIL frameworks to address the few-shot weakly-supervised WSI classification (FSWC) problem by incorporating domain knowledge via text prompts (Figure 1(B)). Multi-scale VLM-based MIL methods [48, 22, 41] further enhance single-scale models by using scale-specific prompts that capture WSI multi-scale information to better represent hierarchical tissue structures.
Although recent multiscale VLM-based MIL methods [48, 22, 41] have made impressive progress, effectively transferring VLM knowledge to MIL remains a serious challenge due to limited modeling of the complex hierarchical structure of WSIs and insufficient integration of multiple modalities. Specifically, as illustrated in Figure 1(C), existing methods face two key limitations. (1) Insufficient modeling of hierarchical interactions within the same modalities. Existing models process visual and textual features independently at each scale and combine them through simple summation or averaging at the final prediction stage. This naïve fusion fails to capture hierarchical relationships across scales within each modality. In the visual domain, it overlooks the semantic progression from coarse tissue-level patterns to fine-grained cellular morphology. Also, in the textual domain, it fails to represent the transition from general morphological descriptions to specific structural details, limiting the model’s ability to leverage hierarchical semantics effectively. (2) Inadequate alignment between modalities on the same scale. Existing models do not fully explore interactions between modalities when constructing task-specific knowledge, often relying on simpler alternative mechanisms that lack the strong inductive bias [4, 58] needed for fine-grained cross-modal alignment. This limits their ability to effectively integrate visual and textual features.
@@ 2.1 Multiple Instance Learning in CPath
WSI classification is typically formulated as a weakly supervised learning task in the MIL setting, where each slide is treated as a bag of unlabeled patches with supervision provided only at the slide level [14]. Embedding-based MIL approaches [27, 33, 38, 47, 54] are generally more effective than instance-based models [7, 9, 43], as they learn discriminative patch embeddings for aggregation [6]. Early methods rely on non-parametric aggregators (e.g., mean, max), while attention-based techniques such as ABMIL [27], DSMIL [33], and CLAM [38] introduce mechanisms to weigh patch relevance. TransMIL [47] captures spatial dependencies via self-attention, while DTFD-MIL [54] employs a double-tier distillation framework with pseudo-bag supervision. Graph-based models [57, 17, 24, 34] improve contextual modeling by constructing structured graphs that capture spatial and relational dependencies between instances. Existing methods remain ineffective for FSWC due to their reliance on large labeled datasets and visual-only features, which make them sensitive to staining variability [36, 11] and domain shifts [37, 48], highlighting the need for a data-efficient approach that leverages prior domain knowledge for robust learning under limited supervision.
@@ 2.2 Vision-Language Models in CPath
Vision-language foundation models such as CLIP [46], BLIP [35], and Flamingo [2] learn joint image-text embeddings through contrastive learning and enable zero- and few-shot transfer via prompting [30, 55, 49, 12]. In CPath, PLIP [25], QuiltNet [26], and CONCH [39] adapt these models using large-scale pathology image-text datasets to enhance robustness in patch-level tasks. Motivated by their few-shot capabilities, recent works extend VLMs to MIL settings for WSI classification. However, adaptation methods developed for natural images, such as CoOp [59], CLIP-Adapter [15], and HeGraphAdapter [56], overlook the gigapixel scale, hierarchical structure, and fine-grained semantics of WSIs, limiting their effectiveness in weakly supervised scenarios. To address this issue, several studies integrate VLMs into MIL frameworks [44, 45, 48, 22, 19]. TOP [44] and FOCUS [19] adopt prompt-based supervision and multi-stage compression, respectively, but are limited to single-scale inputs and cannot capture multi-scale context. Multi-scale VLM-based MIL methods [48, 22, 41] improve upon these approaches by introducing scale-specific prompts that reflect the hierarchical nature of WSIs, enabling more context-aware and semantically aligned representations. Nevertheless, they still face two key challenges: they fail to model semantic progression across scales within each modality and they lack alignment between visual and textual features on the same scale, weakening semantic grounding and multimodal integration. Unlike previous work, HiVE-MIL explicitly addresses both limitations by modeling hierarchical relationships across scales and aligning visual and textual features within each scale, enabling more robust performance in FSWC.
@@ 3.1 Multi-scale Hierarchical Feature Extraction
Unlike prior work [48, 22, 41] that employs multi-scale text prompts and patch features without modeling explicit hierarchical relationships, our method organizes visual and textual representations into a parent–child hierarchy, where each fine-scale (child) node is explicitly linked to its corresponding coarse-scale (parent) node. This hierarchical design reflects the diagnostic workflow of pathologists and captures the intrinsic multi-scale structure of WSIs.
@@
This yields hierarchical textual descriptions: $\textit{Low-scale Text}_{o}$ and $\textit{High-scale Text}_{o,k}$, where each high-scale (child) text is associated with its corresponding low-scale (parent) text, enabling semantically consistent hierarchical alignment. Following CoOp [59], each text is prepended with $L$ learnable tokens $v^{(*)}\in\mathbb{R}^{L\times D}$, resulting in prompt embeddings:
@@ 3.4 Hierarchical Heterogeneous Graph Learning
Intra-scale Aggregator. To capture intra-scale relationships between patch and text nodes on the same scale, we apply a relation-specific GraphSAGE [21] operator $\text{SAGE}^{(r)}(v)$ for each edge type $r\in\mathcal{R}^{\text{intra}}$ ($z_{n}^{(l)}\!\leftrightarrow\!e_{o}^{(l)}$, $z_{r}^{(h)}\!\leftrightarrow\!e_{s}^{(h)}$). The intra-scale representation is then computed by averaging the outputs over all such relations: $h_{v}^{\text{intra}}=\text{MEAN}(\{\text{SAGE}^{(r)}(v)\mid r\in\mathcal{R}^{\text{intra}}\})$. Initial node features $h_{v}$ are modality-specific embeddings.
@@ 4.1 Experimental Settings
Datasets. We utilize three publicly available WSI datasets: TCGA-NSCLC (lung), TCGA-BRCA (breast), and TCGA-RCC (kidney), obtained from The Cancer Genome Atlas (TCGA). [^1] https://portal.gdc.cancer.gov/ Following [48], each dataset is split into training, validation, and test sets using a fixed 4:3:3 ratio. For the few-shot setting, we randomly sample 4, 8, and 16 WSIs per class from the training set. For detailed dataset statistics and preprocessing steps, please refer to the Appendix A.1.
VLM and Baselines. We evaluate using three pathology vision-language foundation models: PLIP [25], QuiltNet [26], and the recent CONCH [39]. Baselines include: (1) Pooling-based (max, mean); (2) Traditional MIL-based (ABMIL [27], DSMIL [33], CLAM-SB/MB [38], TransMIL [47], DTFD-MIL (AFS) [54], WiKG [34]); (3) Single-scale VLM-based MIL method (FOCUS [19]); (4) Multi-scale VLM-based MIL methods (ViLa-MIL [48], MSCPT [22]). All single-scale models including the pooling-based use $20\times$ patches, whereas multi-scale models use $5\times$ and $20\times$. Please refer to the Appendix I for more details.
Implementation Details. HiVE-MIL operates on $5\times$ and $20\times$ patches, using GPT-4o [1] to generate $O=4$ coarse-level texts and $K=3$ fine-level substructures per class. We use $L=16$ learnable context tokens (Eq. 1) and apply the TGDF threshold $\alpha=0.5$ (Eqs. 23). The HHG consists of two layers and a 2-head in MSA. HTCL is used with $\lambda=0.5$ (Eq. 9). We train using Adam optimizer [31] (learning rate: $1\text{e}{-4}$, weight decay: $1\text{e}{-5}$), batch size 1, for up to 50 epochs with early stopping (patience 10). All experiments are run using PyTorch [42] on a workstation with two NVIDIA RTX A100 GPUs. Please refer to the Appendix H for additional implementation details.
@@
PLIP [25] 208K Pathology Image-Text Pairs Max Pooling 55.00 ±3.88 57.33 ±4.63 53.96 ±4.86 57.29 ±3.23 62.33 ±2.94 53.68 ±6.91 66.82 ±6.94 80.58 ±6.68 61.38 ±8.68
@@
ABMIL [27] 70.64 ±2.98 78.44 ±3.63 70.37 ±3.09 65.83 ±5.33 72.87 ±7.88 65.29 ±5.78 80.00 ±3.71 93.01 ±1.53 77.95 ±3.43
DSMIL [33] 72.63 ±3.88 79.88 ±4.60 72.48 ±3.96 71.38 ±3.20 77.55 ±1.62 71.04 ±3.40 86.74 ±1.23 96.44 ±0.63 84.63 ±1.51
CLAM-SB [38] 75.96 ±2.60 83.79 ±3.21 75.94 ±2.61 71.75 ±3.57 80.00 ±2.59 71.49 ±3.60 85.98 ±1.51 96.22 ±0.48 83.35 ±1.54
CLAM-MB [38] 73.46 ±3.15 82.13 ±3.41 73.42 ±3.13 72.50 ±2.92 78.39 ±2.95 72.20 ±2.87 86.97 ±1.03 96.53 ±0.78 84.92 ±1.03
TransMIL [47] 73.21 ±3.02 81.44 ±2.75 72.98 ±2.95 72.08 ±3.32 79.47 ±3.71 71.94 ±3.34 87.05 ±1.52 96.51 ±0.56 84.96 ±1.32
DTFD-MIL [54] 72.95 ±3.40 79.79 ±4.65 72.91 ±3.39 71.25 ±2.68 78.91 ±3.16 70.86 ±2.76 86.74 ±0.79 95.94 ±0.62 84.86 ±1.45
WiKG [34] 67.89 ±3.66 75.54 ±4.05 67.51 ±3.62 67.71 ±2.19 74.92 ±4.16 67.15 ±2.42 83.07 ±0.89 94.34 ±0.76 80.32 ±1.40
ViLa-MIL [48] 74.17 ±1.01 80.63 ±2.37 73.90 ±1.15 71.04 ±6.92 78.42 ±5.86 70.56 ±6.98 85.06 ±2.13 95.53 ±0.97 82.51 ±2.30
MSCPT [22] 76.86 ±1.85 84.93 ±1.59 76.82 ±1.89 72.71 ±2.90 79.78 ±4.14 72.58 ±2.81 86.21 ±0.54 95.84 ±0.45 84.20 ±0.81
FOCUS [19] 71.73 ±5.52 78.21 ±5.93 71.65 ±5.51 71.66 ±5.60 78.19 ±4.51 71.36 ±5.69 87.82 ±1.69 96.73 ±0.70 85.54 ±1.87
@@
QuiltNet [26] 1M Pathology Image-Text Pairs Max Pooling 53.59 ±3.66 57.24 ±5.97 51.36 ±5.39 55.83 ±4.04 56.64 ±4.36 53.75 ±4.57 68.28 ±6.77 81.33 ±7.72 61.31 ±10.73
@@
ABMIL [27] 67.31 ±4.64 75.18 ±5.13 66.81 ±5.22 68.96 ±4.86 76.84 ±4.27 68.42 ±5.45 88.89 ±1.71 96.86 ±0.84 87.11 ±2.44
DSMIL [33] 72.76 ±3.42 78.99 ±3.90 72.53 ±3.41 72.29 ±3.64 79.46 ±2.20 72.06 ±3.54 88.89 ±1.71 96.86 ±0.01 87.11 ±2.44
CLAM-SB [38] 72.82 ±2.68 79.47 ±2.93 72.58 ±2.74 71.46 ±3.82 80.09 ±1.80 71.24 ±4.00 88.66 ±2.17 97.58 ±0.01 87.00 ±2.98
CLAM-MB [38] 73.27 ±3.56 80.53 ±3.76 73.25 ±3.55 72.29 ±2.43 78.42 ±2.75 72.24 ±2.47 88.74 ±1.62 97.34 ±0.01 86.83 ±2.50
TransMIL [47] 71.60 ±4.62 78.59 ±4.86 71.21 ±5.00 71.67 ±3.75 78.77 ±2.92 71.56 ±3.73 86.97 ±1.83 96.71 ±0.01 85.01 ±2.65
DTFD-MIL [54] 70.51 ±5.77 77.38 ±5.26 70.33 ±5.89 72.71 ±2.02 79.28 ±1.81 72.66 ±1.99 88.66 ±1.65 96.74 ±0.71 87.06 ±1.99
WiKG-MIL [34] 68.20 ±3.47 75.08 ±4.66 67.98 ±3.56 68.75 ±3.16 75.51 ±2.16 68.59 ±3.07 83.99 ±1.70 95.13 ±0.70 81.54 ±3.14
ViLa-MIL [48] 73.27 ±5.54 80.82 ±6.41 73.24 ±5.52 72.50 ±3.93 77.67 ±3.12 72.35 ±3.92 84.60 ±1.04 95.67 ±0.70 81.42 ±1.04
MSCPT [22] 76.15 ±3.83 84.06 ±3.02 76.13 ±3.82 72.08 ±5.16 78.59 ±4.21 71.82 ±5.21 87.20 ±1.90 96.89 ±0.87 85.33 ±2.41
FOCUS [19] 69.04 ±3.54 74.64 ±4.29 69.00 ±3.56 68.75 ±4.42 75.66 ±2.86 68.47 ±4.70 89.12 ±1.23 97.13 ±0.46 87.43 ±1.68
@@
CONCH [39] 1.17M Pathology Image-Text Pairs Max Pooling 78.85 ±1.78 87.43 ±1.69 78.82 ±1.77 71.25 ±2.99 78.46 ±4.53 70.91 ±3.14 80.15 ±4.86 91.95 ±2.76 78.11 ±4.60

### v5

@@ 1 Introduction
Whole slide image (WSI) classification is a central task in computational pathology (CPath), enabling cancer diagnosis, subtyping, and prognosis prediction [WSI_Importance_1, WSI_Importance_2, WSI_Importance_3]. With gigapixel resolution, WSIs contain detailed spatial information ranging from coarse tissue structures to fine-grained cellular morphology [WSI_Explanation], which is crucial for accurate interpretation in these diagnostic tasks. To handle their large size and the absence of fine-grained annotations, multiple instance learning (MIL) is widely adopted [MIL, MIL_Histo]. In MIL, each WSI is treated as a bag of instances (patches), and a slide-level label is predicted through a feature aggregator with supervision only at the slide level (Figure 1(A)). However, traditional MIL models [ABMIL, DSMIL, CLAM, TransMIL, DTFD-MIL, WiKG] face key challenges in real-world clinical settings [Real_World_Settings]: (1) they rely on large labeled datasets, which are difficult to obtain due to privacy concerns [Privacy] and the rarity of certain diseases [Rare_Disease_1], making them ineffective in few-shot scenarios where labeled WSIs are scarce [TOP, ViLa-MIL]; (2) they use only visual features, making them highly sensitive to staining variability [Staining_Variability_1, Staining_Variability_2] and domain shifts [IBMIL, ViLa-MIL]. These challenges call for a data-efficient approach that leverages prior domain knowledge for robust learning under limited supervision.
@@
Vision-language models (VLMs) such as CLIP [CLIP], BLIP [BLIP], and Flamingo [Flamingo] have demonstrated strong zero- and few-shot transfer performance by learning joint image-text embeddings through contrastive learning [Zero_Shot_Prompting_1, Zero_Shot_Prompting_2, Few_Shot_Prompting_1, Few_Shot_Prompting_2]. However, their training in natural images and generic captions limits their effectiveness in CPath, where domain-specific semantics are critical. To address this issue, domain-adapted VLMs such as PLIP [PLIP], QuiltNet [QuiltNet], and CONCH [CONCH] leverage large-scale pathology patch-text datasets and improve robustness and generalization at the patch level. Motivated by these capabilities, recent efforts [TOP, PEMP, FOCUS] have integrated them into MIL frameworks to address the few-shot weakly-supervised WSI classification (FSWC) problem by incorporating domain knowledge via text prompts (Figure 1(B)). Multi-scale VLM-based MIL methods [ViLa-MIL, MSCPT, MGPATH] further enhance single-scale models by using scale-specific prompts that capture WSI multi-scale information to better represent hierarchical tissue structures.
Although recent multiscale VLM-based MIL methods [ViLa-MIL, MSCPT, MGPATH] have made impressive progress, effectively transferring VLM knowledge to MIL remains a serious challenge due to limited modeling of the complex hierarchical structure of WSIs and insufficient integration of multiple modalities. Specifically, as illustrated in Figure 1(C), existing methods face two key limitations. (1) Insufficient modeling of hierarchical interactions within the same modalities. Existing models process visual and textual features independently at each scale and combine them through simple summation or averaging at the final prediction stage. This naïve fusion fails to capture hierarchical relationships across scales within each modality. In the visual domain, it overlooks the semantic progression from coarse tissue-level patterns to fine-grained cellular morphology. Also, in the textual domain, it fails to represent the transition from general morphological descriptions to specific structural details, limiting the model’s ability to leverage hierarchical semantics effectively. (2) Inadequate alignment between modalities on the same scale. Existing models do not fully explore interactions between modalities when constructing task-specific knowledge, often relying on simpler alternative mechanisms that lack the strong inductive bias [Inductive_Bias_1, Inductive_Bias_2] needed for fine-grained cross-modal alignment. This limits their ability to effectively integrate visual and textual features.
@@ 2.1 Multiple Instance Learning in CPath
WSI classification is typically formulated as a weakly supervised learning task in the MIL setting, where each slide is treated as a bag of unlabeled patches with supervision provided only at the slide level [MIL_Histo_Review]. Embedding-based MIL approaches [ABMIL, DSMIL, CLAM, TransMIL, DTFD-MIL] are generally more effective than instance-based models [Instance_MIL_1, Instance_MIL_2, Instance_MIL_3], as they learn discriminative patch embeddings for aggregation [Embedding_MIL_Effectiveness]. Early methods rely on non-parametric aggregators (e.g., mean, max), while attention-based techniques such as ABMIL [ABMIL], DSMIL [DSMIL], and CLAM [CLAM] introduce mechanisms to weigh patch relevance. TransMIL [TransMIL] captures spatial dependencies via self-attention, while DTFD-MIL [DTFD-MIL] employs a double-tier distillation framework with pseudo-bag supervision. Graph-based models [Graph_MIL_1, Graph_MIL_2, Graph_MIL_3, WiKG] improve contextual modeling by constructing structured graphs that capture spatial and relational dependencies between instances. Existing methods remain ineffective for FSWC due to their reliance on large labeled datasets and visual-only features, which make them sensitive to staining variability [Staining_Variability_1, Staining_Variability_2] and domain shifts [IBMIL, ViLa-MIL], highlighting the need for a data-efficient approach that leverages prior domain knowledge for robust learning under limited supervision.
@@ 2.2 Vision-Language Models in CPath
Vision-language foundation models such as CLIP [CLIP], BLIP [BLIP], and Flamingo [Flamingo] learn joint image-text embeddings through contrastive learning and enable zero- and few-shot transfer via prompting [Zero_Shot_Prompting_1, Zero_Shot_Prompting_2, Few_Shot_Prompting_1, Few_Shot_Prompting_2]. In CPath, PLIP [PLIP], QuiltNet [QuiltNet], and CONCH [CONCH] adapt these models using large-scale pathology image-text datasets to enhance robustness in patch-level tasks. Motivated by their few-shot capabilities, recent works extend VLMs to MIL settings for WSI classification. However, adaptation methods developed for natural images, such as CoOp [CoOp], CLIP-Adapter [CLIP-Adapter], and HeGraphAdapter [HeGraphAdapter], overlook the gigapixel scale, hierarchical structure, and fine-grained semantics of WSIs, limiting their effectiveness in weakly supervised scenarios. To address this issue, several studies integrate VLMs into MIL frameworks [TOP, PEMP, ViLa-MIL, MSCPT, FOCUS]. TOP [TOP] and FOCUS [FOCUS] adopt prompt-based supervision and multi-stage compression, respectively, but are limited to single-scale inputs and cannot capture multi-scale context. Multi-scale VLM-based MIL methods [ViLa-MIL, MSCPT, MGPATH] improve upon these approaches by introducing scale-specific prompts that reflect the hierarchical nature of WSIs, enabling more context-aware and semantically aligned representations. Nevertheless, they still face two key challenges: they fail to model semantic progression across scales within each modality and they lack alignment between visual and textual features on the same scale, weakening semantic grounding and multimodal integration. Unlike previous work, HiVE-MIL explicitly addresses both limitations by modeling hierarchical relationships across scales and aligning visual and textual features within each scale, enabling more robust performance in FSWC.
@@ 3.1 Multi-scale Hierarchical Feature Extraction
Unlike prior work [ViLa-MIL, MSCPT, MGPATH] that employs multi-scale text prompts and patch features without modeling explicit hierarchical relationships, our method organizes visual and textual representations into a parent–child hierarchy, where each fine-scale (child) node is explicitly linked to its corresponding coarse-scale (parent) node. This hierarchical design reflects the diagnostic workflow of pathologists and captures the intrinsic multi-scale structure of WSIs.
@@
This yields hierarchical textual descriptions: $\textit{Low-scale Text}_{o}$ and $\textit{High-scale Text}_{o,k}$, where each high-scale (child) text is associated with its corresponding low-scale (parent) text, enabling semantically consistent hierarchical alignment. Following CoOp [CoOp], each text is prepended with $L$ learnable tokens $v^{(*)}\in\mathbb{R}^{L\times D}$, resulting in prompt embeddings:
@@ 3.4 Hierarchical Heterogeneous Graph Learning
Intra-scale Aggregator. To capture intra-scale relationships between patch and text nodes on the same scale, we apply a relation-specific GraphSAGE [GraphSAGE] operator $\text{SAGE}^{(r)}(v)$ for each edge type $r\in\mathcal{R}^{\text{intra}}$ ($z_{n}^{(l)}\!\leftrightarrow\!e_{o}^{(l)}$, $z_{r}^{(h)}\!\leftrightarrow\!e_{s}^{(h)}$). The intra-scale representation is then computed by averaging the outputs over all such relations: $h_{v}^{\text{intra}}=\text{MEAN}(\{\text{SAGE}^{(r)}(v)\mid r\in\mathcal{R}^{\text{intra}}\})$. Initial node features $h_{v}$ are modality-specific embeddings.
@@ 4.1 Experimental Settings
Datasets. We utilize three publicly available WSI datasets: TCGA-NSCLC (lung), TCGA-BRCA (breast), and TCGA-RCC (kidney), obtained from The Cancer Genome Atlas (TCGA). [^1] https://portal.gdc.cancer.gov/ Following [ViLa-MIL], each dataset is split into training, validation, and test sets using a fixed 4:3:3 ratio. For the few-shot setting, we randomly sample 4, 8, and 16 WSIs per class from the training set. For detailed dataset statistics and preprocessing steps, please refer to the Appendix A.1.
VLM and Baselines. We evaluate using three pathology vision-language foundation models: PLIP [PLIP], QuiltNet [QuiltNet], and the recent CONCH [CONCH]. Baselines include: (1) Pooling-based (max, mean); (2) Traditional MIL-based (ABMIL [ABMIL], DSMIL [DSMIL], CLAM-SB/MB [CLAM], TransMIL [TransMIL], DTFD-MIL (AFS) [DTFD-MIL], WiKG [WiKG]); (3) Single-scale VLM-based MIL method (FOCUS [FOCUS]); (4) Multi-scale VLM-based MIL methods (ViLa-MIL [ViLa-MIL], MSCPT [MSCPT]). All single-scale models including the pooling-based use $20\times$ patches, whereas multi-scale models use $5\times$ and $20\times$. Please refer to the Appendix I for more details.
Implementation Details. HiVE-MIL operates on $5\times$ and $20\times$ patches, using GPT-4o [GPT-4o] to generate $O=4$ coarse-level texts and $K=3$ fine-level substructures per class. We use $L=16$ learnable context tokens (Eq. 1) and apply the TGDF threshold $\alpha=0.5$ (Eqs. 23). The HHG consists of two layers and a 2-head in MSA. HTCL is used with $\lambda=0.5$ (Eq. 9). We train using Adam optimizer [Adam] (learning rate: $1\text{e}{-4}$, weight decay: $1\text{e}{-5}$), batch size 1, for up to 50 epochs with early stopping (patience 10). All experiments are run using PyTorch [PyTorch] on a workstation with two NVIDIA RTX A100 GPUs. Please refer to the Appendix H for additional implementation details.
@@
PLIP [PLIP] 208K Pathology Image-Text Pairs Max Pooling 55.00 ±3.88 57.33 ±4.63 53.96 ±4.86 57.29 ±3.23 62.33 ±2.94 53.68 ±6.91 66.82 ±6.94 80.58 ±6.68 61.38 ±8.68
@@
ABMIL [ABMIL] 70.64 ±2.98 78.44 ±3.63 70.37 ±3.09 65.83 ±5.33 72.87 ±7.88 65.29 ±5.78 80.00 ±3.71 93.01 ±1.53 77.95 ±3.43
DSMIL [DSMIL] 72.63 ±3.88 79.88 ±4.60 72.48 ±3.96 71.38 ±3.20 77.55 ±1.62 71.04 ±3.40 86.74 ±1.23 96.44 ±0.63 84.63 ±1.51
CLAM-SB [CLAM] 75.96 ±2.60 83.79 ±3.21 75.94 ±2.61 71.75 ±3.57 80.00 ±2.59 71.49 ±3.60 85.98 ±1.51 96.22 ±0.48 83.35 ±1.54
CLAM-MB [CLAM] 73.46 ±3.15 82.13 ±3.41 73.42 ±3.13 72.50 ±2.92 78.39 ±2.95 72.20 ±2.87 86.97 ±1.03 96.53 ±0.78 84.92 ±1.03
TransMIL [TransMIL] 73.21 ±3.02 81.44 ±2.75 72.98 ±2.95 72.08 ±3.32 79.47 ±3.71 71.94 ±3.34 87.05 ±1.52 96.51 ±0.56 84.96 ±1.32
DTFD-MIL [DTFD-MIL] 72.95 ±3.40 79.79 ±4.65 72.91 ±3.39 71.25 ±2.68 78.91 ±3.16 70.86 ±2.76 86.74 ±0.79 95.94 ±0.62 84.86 ±1.45
WiKG [WiKG] 67.89 ±3.66 75.54 ±4.05 67.51 ±3.62 67.71 ±2.19 74.92 ±4.16 67.15 ±2.42 83.07 ±0.89 94.34 ±0.76 80.32 ±1.40
ViLa-MIL [ViLa-MIL] 74.17 ±1.01 80.63 ±2.37 73.90 ±1.15 71.04 ±6.92 78.42 ±5.86 70.56 ±6.98 85.06 ±2.13 95.53 ±0.97 82.51 ±2.30
MSCPT [MSCPT] 76.86 ±1.85 84.93 ±1.59 76.82 ±1.89 72.71 ±2.90 79.78 ±4.14 72.58 ±2.81 86.21 ±0.54 95.84 ±0.45 84.20 ±0.81
FOCUS [FOCUS] 71.73 ±5.52 78.21 ±5.93 71.65 ±5.51 71.66 ±5.60 78.19 ±4.51 71.36 ±5.69 87.82 ±1.69 96.73 ±0.70 85.54 ±1.87
@@
QuiltNet [QuiltNet] 1M Pathology Image-Text Pairs Max Pooling 53.59 ±3.66 57.24 ±5.97 51.36 ±5.39 55.83 ±4.04 56.64 ±4.36 53.75 ±4.57 68.28 ±6.77 81.33 ±7.72 61.31 ±10.73
@@
ABMIL [ABMIL] 67.31 ±4.64 75.18 ±5.13 66.81 ±5.22 68.96 ±4.86 76.84 ±4.27 68.42 ±5.45 88.89 ±1.71 96.86 ±0.84 87.11 ±2.44
DSMIL [DSMIL] 72.76 ±3.42 78.99 ±3.90 72.53 ±3.41 72.29 ±3.64 79.46 ±2.20 72.06 ±3.54 88.89 ±1.71 96.86 ±0.01 87.11 ±2.44
CLAM-SB [CLAM] 72.82 ±2.68 79.47 ±2.93 72.58 ±2.74 71.46 ±3.82 80.09 ±1.80 71.24 ±4.00 88.66 ±2.17 97.58 ±0.01 87.00 ±2.98
CLAM-MB [CLAM] 73.27 ±3.56 80.53 ±3.76 73.25 ±3.55 72.29 ±2.43 78.42 ±2.75 72.24 ±2.47 88.74 ±1.62 97.34 ±0.01 86.83 ±2.50
TransMIL [TransMIL] 71.60 ±4.62 78.59 ±4.86 71.21 ±5.00 71.67 ±3.75 78.77 ±2.92 71.56 ±3.73 86.97 ±1.83 96.71 ±0.01 85.01 ±2.65
DTFD-MIL [DTFD-MIL] 70.51 ±5.77 77.38 ±5.26 70.33 ±5.89 72.71 ±2.02 79.28 ±1.81 72.66 ±1.99 88.66 ±1.65 96.74 ±0.71 87.06 ±1.99
WiKG-MIL [WiKG] 68.20 ±3.47 75.08 ±4.66 67.98 ±3.56 68.75 ±3.16 75.51 ±2.16 68.59 ±3.07 83.99 ±1.70 95.13 ±0.70 81.54 ±3.14
ViLa-MIL [ViLa-MIL] 73.27 ±5.54 80.82 ±6.41 73.24 ±5.52 72.50 ±3.93 77.67 ±3.12 72.35 ±3.92 84.60 ±1.04 95.67 ±0.70 81.42 ±1.04
MSCPT [MSCPT] 76.15 ±3.83 84.06 ±3.02 76.13 ±3.82 72.08 ±5.16 78.59 ±4.21 71.82 ±5.21 87.20 ±1.90 96.89 ±0.87 85.33 ±2.41
FOCUS [FOCUS] 69.04 ±3.54 74.64 ±4.29 69.00 ±3.56 68.75 ±4.42 75.66 ±2.86 68.47 ±4.70 89.12 ±1.23 97.13 ±0.46 87.43 ±1.68
@@
CONCH [CONCH] 1.17M Pathology Image-Text Pairs Max Pooling 78.85 ±1.78 87.43 ±1.69 78.82 ±1.77 71.25 ±2.99 78.46 ±4.53 70.91 ±3.14 80.15 ±4.86 91.95 ±2.76 78.11 ±4.60